OpenStudy (anonymous):

A triangle has a perimeter of 50. If two of the side lengths are equal and the third side is five more than the equal sides, what is the length of the third side?

OpenStudy (karatechopper):

ok! do you like to solve questions in drawing or equations?

OpenStudy (anonymous):

@karatechopper will tell you how to do this..

OpenStudy (ganpat):

say the three side of triangle are a, b and c... a + b + c = 50 and say, a= b and c = a + 5 so r equation becomes, a + a + a + 5 = 50 3a + 5 = 50 3a = 45 a = 15 = b and c = a + 5 = 15 + 5 = 20